Off-The-Books Investment Firm, LLC, has offered you an investment it says will return to you $20,000 in 2 years. To get in, you'll need to make a $10,000 deposit to their receivables account and promise not to tell anyone about it. What is the annual return on this investment?

Answer: 41% however what is the excel calculation/formula used to get 41%?
Answer #1

We know that annual return =  maturity futurevalue present value

Future value = $20000

Present value = $10000

Hence annual return = 2/20000 V 10000

=| |-

=1.414 - 1 = 41.4 = 41% (approx.)